OpenStudy (anonymous):

solve the equation 33 =p-6.71 can someone help me with this? -

OpenStudy (jhannybean):

evaluate for p, \[33=p-6.71\]\[33+6.71 = p-6.71+6.71\]

OpenStudy (anonymous):

i need help figuring out how to find what p means i am very confused...

OpenStudy (jhannybean):

p is a variable, just like x or y,according to this equation

OpenStudy (anonymous):

so do i add 33 to 6.71 to get p?

OpenStudy (kenljw):

If 33=p^-6.71 log(p) = -log(33)/6.71 p=10^-log(33)/6.7)

OpenStudy (jhannybean):

yes you do :)

OpenStudy (jhannybean):

@KenLJW I'm not sure if this person has learned what logarithmic functions are yet..... maybe.

OpenStudy (anonymous):

so that would make p = 39.71

OpenStudy (jhannybean):

:\

OpenStudy (kenljw):

I've assume the question as stated was to simple, the result only requires basic math and minimum rudiments of algebra

OpenStudy (jhannybean):

Some people are beginning to learn math? :P

OpenStudy (kenljw):

On this supposedly College level? or are the Algebra courses here HS level? Is there a course for College algebra, or are they all preparatory?